# Oscillatrons: neural units with time-dependent multifunctionality

### v2

@@ II Oscillatrons are multifunctional in a time-dependent way
Conventional physical models of processes such as bit erasure consider the manipulation of a degree of freedom in an external potential, with work expended Landauer (1961); Sagawa (2012); Zulkowski and DeWeese (2014); Proesmans et al. (2020); Dago and Bellon (2022). Here, energy must be input in order to set the initial state of the unit and impose and maintain the input to the neuron (which sets the potential spring constant). But after setting the initial condition no work is performed: computations are done by the system’s natural dynamics  [^1] Single operations and multiple operations incur the same energy cost, with the caveat that the input $I$ to the neuron must be maintained long enough to observe these operations..
Now consider two binary input degrees of freedom $I_{1,2}=0,1$, and construct the input $I=I_{1}+I_{2}$. $I$ can take three values, 0,1, or 2. In Fig. 1(b) we show the oscillatron output (2) as a function of time $t$ for the three possible values of $I$. With the horizontal decision boundary placed as shown, the neuron can function as any of the elementary logic gates, depending on observation time. If we observe the unit at time $t_{1}$ it functions as an AND gate; at time $t_{2}$, an OR gate; and at time $t_{3}$, an XOR gate. At different times the same unit can also function as the inverted versions of these gates, NAND, XNOR, and NOR, the latter shown at time $t_{4}$.
The equilibrium state of the unit is $S(t\to\infty)=0$, regardless of the value of $I$, and so in equilibrium the unit has no computational ability. In general, we can do time-resolved computation in or out of equilibrium (to do so here we would add a source term to the right-hand-side of Eq. (1)), but doing it out of equilibrium means that we do not need to wait for equilibrium to be attained.
III Training oscillatrons to perform multifunctional computation
Oscillatrons can be trained by gradient descent to achieve distinct tasks at distinct times. To illustrate this point we consider the MNIST data set LeCun et al. (1998), which consists of greyscale images of $70,000$ handwritten digits on a grid of $N=28\times 28$ pixels, each digit belonging to one of ten classes $C\in[0,9]$; see Fig. 2(a). We wish to construct a classifier that discriminates 0s (class $C=0$) from 1s (class $C=1$) at observation time $t_{1}$; discriminates 4s from 5s at time $t_{2}$; and discriminates 8s from 9s at time $t_{3}$. We choose observation times $t_{1}$, $t_{2}$, and $t_{3}$ such that $\Omega_{0}t_{k}=-\arctan(2\Omega_{0}/\gamma)+(k+3)\pi$, where $\Omega_{0}^{2}\equiv\omega_{0}^{2}-\gamma^{2}/4$, corresponding to three of the roots of the function (2) with $I=0$.
@@ III Training oscillatrons to perform multifunctional computation
Here $\alpha$ labels MNIST digits, and the inner sum runs over training-set digits in classes $C_{4k}$ and $C_{4k+1}$. We have defined $s_{\alpha}(t_{k})\equiv{\rm e}^{\gamma t/2}S_{\alpha}(t_{k})$; the exponential scaling cancels the exponential decay of the solution (2), ensuring that each term in (4) is of equal importance. $S_{\alpha}(t_{k})$ denotes the value of (2) at time $t_{k}$ when shown MNIST digit $\alpha$. We consider $N$ connections between digit and neuron: when shown digit $\alpha$, the neuron input is $I_{\alpha}=\sum_{i=1}^{N}\theta_{i}I_{i,\alpha}$, where $I_{i,\alpha}$ is the value of the $i^{\rm th}$ pixel of MNIST digit $\alpha$ and the $N$ quantities $\theta_{i}$ are adjustable weights. In what follows we use the notation $\Omega_{\alpha}^{2}\equiv\omega_{0}^{2}-\gamma^{2}/4+I_{\alpha}$. The target activity in (4) is $\bar{s}_{\alpha}=\mp 1/2$, where the negative or positive sign is chosen if digit $\alpha$ is a member of an even-numbered class $(0,4,8)$ or an odd-numbered class $(1,5,9)$, respectively.
To train the oscillatron to solve this multi-time classification task we adjust its input weights by gradient descent on $\phi$. Weights are initially chosen randomly, $\theta_{i}\sim{\mathcal{N}}(0,10^{-4})$; we then iterate the equation $\theta_{i}\to\theta_{i}-\alpha_{0}\partial\phi/\partial\theta_{i}$ for all $N$ weights $\theta_{i}$. Here $\alpha_{0}$ is the learning rate and
@@
Results are shown in Fig. 2(b). We plot time traces of oscillatron activity when shown 10 test-set digits (unseen during training) in each class. Classes are indicated by color. The neuron responds differently to each class, and can (imperfectly) distinguish the required classes at the required times: on the right of the panel we show histograms of neuron activity, using all digits of that class in the $10,000$-digit MNIST test set, at the designated observation times.
As with standard neural units, collections of oscillatrons are more expressive than individual oscillatrons. Multiple oscillator neurons can have a total activity that is not periodic in time, even if they do not interact. In Fig. 2(c) we show the analog of panel (b) for a collection of 4 noninteracting oscillatrons. Each has the same parameters $\omega_{0}$ and $\gamma$ as the single oscillatron of panel (b), and each has $N$ input weights. We impose the loss function $\phi$, with $S_{\alpha}(t)$ replaced by $\frac{1}{4}\sum_{i=1}^{4}S^{(i)}_{\alpha}(t)$. Here $S^{(i)}_{\alpha}(t)$ is the output of oscillatron $i$ at time $t$ when shown MNIST digit $\alpha$. As before, we adjust the input weights by gradient descent on the loss.
The time traces and histograms of Fig. 2(c) show the trained neural gas to be more expressive than the trained individual neuron, better distinguishing the required classes at the designated observation time (the loss values for the individual neuron and the gas are shown in Fig. 2(d)). In panel (e) we show histograms of neural activity at a time $t_{3}$ for which the neural gas is trained to discriminate 8s and 9s but not 0s and 1s; there, it cannot distinguish 0s and 1s.
@@ IV Conclusions
Several branches of computing use the natural evolution of a physical system to do calculations. We have shown that the dynamics of an underdamped harmonic oscillator can perform multifunctional computation, with the same physical system able to solve distinct problems at distinct times within a single dynamical trajectory. The idea proposed here, a form of analog computing, is a nonstandard form of oscillator computing: the latter usually focuses on the information contained within the phase of an oscillator, and seeks to identify the ground-state phases of coupled oscillators. Here we have considered the time-resolved amplitude of an oscillator whose inputs influence its frequency, which we interpret as a time-dependent neural unit or oscillatron (by comparison with perceptron). The activity of the oscillatron at fixed time is a nonmonotonic function of the input, and so it can solve nonlinearly-separable problems such as XOR. The activity of the oscillatron at fixed input is a nonmonotonic function of time, and so it is is multifunctional in a temporal sense, able to carry out distinct nonlinear computations at distinct times within the same dynamical trajectory. We have shown that oscillatrons can do all one-bit operations and function as all of the elementary logic gates, depending on observation time, and can be trained by gradient descent to perform distinct classification tasks at distinct times.
@@
A single oscillatron can do nonlinear computation in a multifunctional way, and is a natural building block for neural networks and other devices. Neural networks can be built from interacting oscillators just as they are built from standard artificial neurons Atkins (1991); Effenberger et al. (2022). Oscillator neural nets designed for multifunctional computation would have considerable expressiveness. For instance, connecting $N$ oscillatrons $S_{i}(t)$ with springs or Ising-like potentials $U_{ij}=k_{ij}S_{i}(t)S_{j}(t)$ produces a neural network in which, for fixed input data, the activity of each oscillatron is the superposition of $N$ normal modes, admitting neural waves for large $N$ French (1971); Landau and Lifshitz (1976) (neural oscillations and waves are common in the brain Kandel et al. (2000)). Inter-neuron couplings higher order than bilinear would allow for the possibility of chaotic behavior, resulting in stochastic behavior even in the absence of external noise; this would permit probabilistic and generative computation.
The idea described here does not require thermal noise but could be applied in its presence: in fields such as thermodynamic computing, thermal noise provides an important mechanism for driving dynamical evolution and allowing probabilistic computation Aifer et al. (2023). The results of this paper suggest extending the idea of thermodynamic computing to consider the time-resolved dynamics of devices that already exist, such as networks of interacting oscillators printed on circuit boards Wang and Roychowdhury (2019); Chou et al. (2019); Melanson et al. (2023).
@@
Time-resolved multifunctional computing provides a way of carrying out multiple nonlinear computations within a single dynamical trajectory of a device. This idea could help reduce the number of parameters or the size of devices needed to do computation, with the natural time evolution of a device giving us, in effect, multiple computations for the price of one.
V Acknowledgments.
@@
I thank Isaac Tamblyn for comments on the paper. This work was done at the Molecular Foundry at Lawrence Berkeley National Laboratory, supported by the Office of Basic Energy Sciences of the U.S. Department of Energy under Contract No. DE-AC02–05CH11231.

### v3

@@ III Training oscillatrons by gradient descent to perform multifunctional computation
In this section we show how an oscillator can be trained by gradient descent to achieve distinct tasks at distinct times. We use the oscillator (2) with intrinsic parameters $\omega_{0}=1$ and $\gamma=1/10$. We consider the MNIST data set LeCun et al. (1998), which consists of greyscale images of $70,000$ handwritten digits on a grid of $N=28\times 28$ pixels, each digit belonging to one of ten classes $C\in[0,9]$; see Fig. 4(a). We wish to construct a classifier that discriminates 0s (class $C=0$) from 1s (class $C=1$) at observation time $t_{1}$; discriminates 4s from 5s at time $t_{2}$; and discriminates 8s from 9s at time $t_{3}$. We choose observation times $t_{1}$, $t_{2}$, and $t_{3}$ such that $\Omega_{0}t_{k}=-\arctan(2\Omega_{0}/\gamma)+(k+3)\pi$, where $\Omega_{0}^{2}\equiv\omega_{0}^{2}-\gamma^{2}/4$, corresponding to three of the roots of the function (3) with $I=0$.
@@
Here $\alpha$ labels MNIST digits, and the inner sum runs over training-set digits in classes $C_{4k}$ and $C_{4k+1}$. We have defined $s_{\alpha}(t_{k})\equiv{\rm e}^{\gamma t/2}S_{\alpha}(t_{k})$; the exponential scaling cancels the exponential decay of the solution (3), ensuring that each term in (8) is of equal importance. $S_{\alpha}(t_{k})$ denotes the value of (3) at time $t_{k}$ when shown MNIST digit $\alpha$. The target activity in (8) is $\bar{s}_{\alpha}=\mp 1/2$, where the negative or positive sign is chosen if digit $\alpha$ is a member of an even-numbered class $(0,4,8)$ or an odd-numbered class $(1,5,9)$, respectively. The outer sum runs over the three observation times.
We consider $N$ connections between digit and neuron: when shown digit $\alpha$, the neuron input is $I_{\alpha}=\sum_{i=1}^{N}\theta_{i}I_{i,\alpha}$. Here $I_{i,\alpha}$ is the value of the $i^{\rm th}$ pixel of MNIST digit $\alpha$, and the $N$ quantities $\theta_{i}$ are adjustable weights. In what follows we use the notation $\Omega_{\alpha}^{2}\equiv\omega_{0}^{2}-\gamma^{2}/4+I_{\alpha}$.
To train an oscillator to solve this multi-time classification task we adjust its input weights by gradient descent on $\phi$. Weights are initially chosen randomly, $\theta_{i}\sim{\mathcal{N}}(0,10^{-4})$; we then iterate the equation
$\theta_{i}\to\theta_{i}-\alpha_{0}\frac{\partial\phi}{\partial\theta_{i}}$ (9)
for all $N$ weights $\theta_{i}$. Here $\alpha_{0}$ is the learning rate and
@@
Results are shown in Fig. 4(b). We plot time traces of oscillator activity when shown 10 test-set digits (unseen during training) in each class. Classes are indicated by color. The oscillator responds differently to each class, and can (imperfectly) distinguish the required classes at the required times: on the right of the panel we show histograms of oscillator activity, using all digits of the indicated class in the $10,000$-digit MNIST test set, at the designated observation times.
Note that training is done on the computer, the idea being that the trained weights can be implemented in hardware. For instance, if currents are used to represent input pixel values, then resistors can be used to encode the weights. This process is distinct to the process in which a physical system adapts to applied forces in order to learn to perform a task Stern et al. (2021).
As with standard neural units, collections of oscillatrons are more expressive than individual units. Multiple oscillators can have a total activity that is not periodic in time, even if they do not interact. In Fig. 4(c) we show the analog of panel (b) for a collection of 4 noninteracting oscillator units (sometimes called a neural gas). Each has the same parameters $\omega_{0}$ and $\gamma$ as the single oscillator of panel (b), and each has $N$ input weights. We impose the loss function $\phi$, with $S_{\alpha}(t)$ replaced by $\frac{1}{4}\sum_{i=1}^{4}S^{(i)}_{\alpha}(t)$. Here $S^{(i)}_{\alpha}(t)$ is the output of unit $i$ at time $t$ when shown MNIST digit $\alpha$. As before, we adjust the input weights by gradient descent on the loss.
The time traces and histograms of Fig. 4(c) show the trained oscillator gas to be more expressive than the trained individual oscillator, better distinguishing the required classes at the designated observation time. The loss values for the individual oscillator and the gas are shown in Fig. 4(d). In panel (e) we show histograms of oscillator activity at a time $t_{3}$ for which the oscillator gas is trained to discriminate 8s and 9s but not 0s and 1s; there, it cannot distinguish 0s and 1s.
@@ IV Conclusions
We have shown that the dynamics of an underdamped harmonic oscillator can perform multifunctional computation, with the same physical system able to solve distinct problems at distinct times within a single dynamical trajectory. This concept can be realized in hardware – (3) is the output of the physical dynamics (2) – and corresponds to a nonstandard form of oscillator computing: the latter usually focuses on the information contained within the phase of an oscillator, and seeks to identify the ground-state phases of coupled oscillators. Here we have considered the time-resolved amplitude of an oscillator whose inputs influence its frequency (or that is observed on a time that is a function of the input), which we interpret as a time-dependent neural unit or oscillatron, by comparison with perceptron. The activity of the oscillatron at fixed time is a nonmonotonic function of the input, and so it can solve nonlinearly-separable problems such as XOR. The activity of the oscillatron at fixed input is a nonmonotonic function of time, and so it is is multifunctional in a temporal sense, able to carry out distinct nonlinear computations at distinct times within the same dynamical trajectory.
We have shown that oscillatrons can do all one-bit binary operations, function as all of the elementary logic gates, perform binary arithmetic, and convert an arbitrary number of analog values to a binary representation. We have also shown that oscillatrons can be trained by gradient descent to perform distinct classification tasks at distinct times. To perform multifunctional computation we must observe oscillators at multiple times. Elements that can be used for timing (such as digital-to-time converters, phase-locked loops, or clocked latches) use power, and power expenditure would rise with the precision and number of time points needed Roberts and Ali-Bakhshian (2010); Best (2007). The energy cost of timing is an important consideration to realize multifunctional computation in practice.
@@
A single oscillator can do nonlinear computation in a multifunctional way, and is a natural building block for neural networks and other devices. Neural networks can be built from interacting oscillators just as they are built from standard artificial neurons Atkins (1991); Vodenicarevic et al. (2017); Effenberger et al. (2022). Oscillator neural nets designed for multifunctional computation would have considerable expressiveness. For instance, connecting $N$ oscillators $S_{i}(t)$ with springs or Ising-like potentials $U_{ij}=k_{ij}S_{i}(t)S_{j}(t)$ produces a neural network in which, for fixed input data, the activity of each oscillator is the superposition of $N$ normal modes, admitting neural waves for large $N$ French (1971); Landau and Lifshitz (1976). Neural oscillations and waves are common in the brain Kandel et al. (2000). Inter-neuron couplings higher order than bilinear would allow for the possibility of chaotic behavior, resulting in stochastic behavior even in the absence of external noise; this would permit probabilistic and generative computation.
@@
Computing with time-dependent functionality provides a way of carrying out multiple nonlinear computations within a single dynamical trajectory of a device, its natural time evolution giving us, in effect, multiple computations for the price of one. This idea could help reduce the number of parameters of a neural network or the number of circuit elements required to do specified computations.
@@
Acknowledgments— I thank Isaac Tamblyn for comments on the paper. This work was done at the Molecular Foundry at Lawrence Berkeley National Laboratory, supported by the Office of Basic Energy Sciences of the U.S. Department of Energy under Contract No. DE-AC02–05CH11231.